OpenStudy (anonymous):

given an = 5n find a3

OpenStudy (goten77):

hmm

OpenStudy (anonymous):

if \[a_1 = 5*1\] and if \[a_2 = 5*2\] then \[a_3 = ???\]

OpenStudy (goten77):

an = 5n this could also be also said as a = 5 since its being multiplied and the n on each side can cancel out a=5 now a x 3 = ??? 5 x 3 = 15